OpenStudy (anonymous):

A car's wheels are 28" in diameter. How far (in miles) will the car travel if its wheels revolve 10,000 times without slipping?

OpenStudy (anonymous):

for one revolution of a wheel car travels 2*pi*14 ins 10000 revs = 2*pi*14* 10,000 ins = ( 2*pi*14* 10,000) / (5280*12) miles = 13.883 miles

OpenStudy (turingtest):

convert miles to inches (12in/ft)(5280ft/mi)=63360in/mi C=(pi)d=28pi=distance of one revolution in inches 28pi(10,000)=280000pi inches convert to miles: 280000/63360=4.419 miles hey I which one of us messed up jimmyrep?

OpenStudy (turingtest):

oh id did (forgot to use pi) 4.419(pi) miles=13.883 miles my bad

OpenStudy (anonymous):

Jimmyrep : where did the 12 come from when multiplying by 5280? and that answer was correct. I have the answers in the back of my book. I just didn't understand how to get it.

OpenStudy (turingtest):

to get inches/mile you must do a conversion: 12(inches/foot)x5280(feet/mile)=63360(inches/mile)

OpenStudy (anonymous):

yes 5280 ft to mile 12 ins to a ft

OpenStudy (turingtest):

 the "feet" cancel